# Rapidly Converging Time-Discounted Ergodicity on Graphs for Active Inspection of Confined Spaces

### v1

@@ Abstract
Ergodic exploration has spawned a lot of interest in mobile robotics due to its ability to design time trajectories that match desired spatial coverage statistics. However, current ergodic approaches are for continuous spaces, which require detailed sensory information at each point and can lead to fractal-like trajectories that cannot be tracked easily. This paper presents a new ergodic approach for graph-based discretization of continuous spaces. It also introduces a new time-discounted ergodicity metric, wherein early visitations of information-rich nodes are weighted more than late visitations. A Markov chain synthesized using a convex program is shown to converge more rapidly to time-discounted ergodicity than the traditional fastest mixing Markov chain. The resultant ergodic traversal method is used within a hierarchical framework for active inspection of confined spaces with the goal of detecting anomalies robustly using SLAM-driven Bayesian hypothesis testing. Both simulation and physical experiments on a ground robot show the advantages of this framework over greedy and random exploration methods for left-behind foreign object debris detection in a ballast tank.
@@ I Introduction
There are three main challenges of continuous-space ergodic control for inspection of confined spaces, all of which are addressed in our discrete-space method. First, the continuous space methods tend to yield complex, fractal-like trajectories to satisfy the ergodicity requirements, which are difficult to execute in confined spaces. Moreover, the complex trajectories (i) can potentially destabilize robot localization by introducing large nonlinearities in the inputs and (ii) can cause the robot to inspect local featureless areas for long periods of time. A graph-based discretization of the confined space into relatively large regions allows designing a region-wise ergodic traversal method that works without considering local collision-free motion planning directly. Second, the information field for the continuous-space ergodic method requires a sensor model to consider occlusion at every point in space. In contrast, the information is bulk estimated at each node on the graph (region in the space) in our discrete ergodic control method. Third, the continuous-space ergodic methods cannot model directional traversability constraints in the confined space, such as an unclimbable slope in one direction, or a door that only opens on one side. In contrast, such connectivity constraints are modeled as directed graph edges in our discrete method.
@@
This graph-based discretization enables the development of a hierarchical planning framework, which allows decomposition and modularization of both robot control and estimation to reduce the overall computational complexity of active inspection. We focus on anomaly detection as the inspection goal. Consequently, REMC sits at the top of the hierarchy as the region-level planner, where coarse-grained information on the current uncertainty estimates of region-wise anomalies are fed from the region graph. At the second level, fine-grained information on anomaly uncertainty estimates within the regions are used for inspection waypoint placement. At the bottom level, robot state estimates and an occupancy grid map are provided as inputs by a graph-SLAM algorithm for local waypoint navigation. The uncertainty estimates come from a Bayesian anomaly detector at this level, which recursively updates the posterior of the robot’s local observations corresponding to anomalies based on the signed Mahalanobis distances of the observations from their nearest points on a reference model of the confined space. The overall inspection framework is termed hierarchical ergodic Markov planning (HEMaP). Both simulation and physical experiments are carried out inside a representative ballast tank with FODs as anomalies, using a ground robot equipped with a depth camera and GPU-based on-board processing. The results show that HEMaP achieves high FOD detection rates (averaging 85%), and consistently outperforms greedy and random exploration methods regardless of the allowable inspection time frame.
@@ II-A Ergodic Control: Continuous Versus Discretized Space
There are three main branches of work that attempt to address these challenges. One branch utilizes the Fourier metric as an objective function and expands the method into a trajectory optimization problem, where collision avoidance is handled as an optimization constraint [16, 17, 18, 19]. However, it inherits a similar problem as the Fourier-based metric, where the finite number of frequency coefficients are not able to capture complex obstacle geometry such as structures in the confined space. When modeled as constraints, the complex geometries, together with the robot motion model, introduce high non-convexity, which generates more computational complexity. The second branch of work uses the radial basis function instead of the Fourier basis to adapt to different domain topologies [20, 21, 22]. The basis functions are specific to the domain topology and are pre-computed using prior maps. When unexpected obstacles such as FODs appear, a new set of basis functions may be required. In contrast, new obstacles are handled by a standard local planner in our hierarchical approach. The third line of work changes the ergodic metric to KL-divergence between the robot distribution and the target distribution. As a result, this metric is able to adapt to a wider range of sensor models and has a lower computational complexity than the Fourier method [23]. This line of work is similar to our approach in that both utilize stochasticity to solve the ergodic control problem. However, instead of finding a stochastic policy that is inherently ergodic, we use stochastic optimization to efficiently sample trajectories that solve the finite-horizon optimization problem.
Several works have been done on adapting these three branches of ergodic control methods on various robotics tasks, including target localization [24], multi-link manipulator control [25, 26, 27, 28], and swarm control [29]. In these tasks, the information field is either given or has a well-behaved analytical form. However, information is harder to model in inspection applications, which instead benefit from a graph-based abstraction.
@@ III-A Hierarchical Representation of Workspace
We consider the set of regions in the workspace as an arbitrary coarse graining of the pose space $\mathcal{X}\subset\text{SE}(3)$ into a finite disjoint set =$\mathcal{R}=\{r_{1},r_{2},\cdots,r_{n}\}$, where the space within each region is connected.
@@ IV-A Continuous and Discrete Ergodic System
for all initial conditions $r[0]$. Here, the right-hand side is the space average of $F$ over all elements $r$ in the set $\mathcal{R}$ with respect to a measure $\mu$ over the sigma algebra $\Sigma$. The left-hand side is the (long-term) time average of $F$ evaluated along the sequence $r[k]$ generated by a dynamic map $\ r[k+1]=M(r[k]):\mathcal{R}\rightarrow\mathcal{R}$ with initial state $r[0]$.
The graph-based ergodicity in (IV.2) implies that the relative frequency of the robot visiting each region (i.e., the time average $\hat{\rho}$) is equal to equal to the relative measure of each region on the graph (i.e., the space average $\bar{\rho}$) if the measurable function $F$ in (3) is selected as the indicator function $I$
@@ IV-B Asymptotic Ergodicity
Asymptotic ergodicity can be achieved if the transition probability between regions is determined by any irreducible stochastic matrix $P$ with the stationary distribution equal to the desired distribution. Specifically, consider a stochastic dynamic map $M$ in (IV.2), where the transition probability $P_{j,i}$ of the robot from region $r_{i}$ to $r_{j}$ forms a stochastic matrix $P\in\mathbb{R}^{n\times n}$. Then, asymptotic ergodicity, in the expected sense, is achieved for any stochastic matrix $P$ that is irreducible, i.e., can go from any region to any other region in the graph $\mathcal{G}$ (not necessarily in one move), with the space average $\bar{\rho}$ as the unique stationary distribution, as shown below.
Theorem 1.
@@
Proof.
By the Perron-Frobenius theorem, if the stochastic matrix $P$ is irreducible, then the Cesàro sum of $P$ exists, and it is equal to an $n\times n$ matrix with columns spanned by the stationary distribution $\bar{\rho}$ (which is unique) [42, p.692], i.e.
$\lim_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\left(P^{k}\right)=\bar{\rho}\textbf{1}^{T}=\begin{bmatrix}\vline&\vline&\\ \bar{\rho}&\bar{\rho}&\cdots\\ \vline&\vline&\\ \end{bmatrix}.$ (11)
Therefore, (8) can be rewritten as
$\mathbb{E}[\hat{\rho}]=\left(\lim_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}P^{k}\right)\rho_{0}=\begin{bmatrix}\vline&\vline&\\ \bar{\rho}&\bar{\rho}&\cdots\\ \vline&\vline&\\ \end{bmatrix}\rho_{0}=\bar{\rho}$ (12)
since $\rho_{0}$ terms sum to one (as it is a distribution), leading to the theorem’s result. ∎
@@ IV-C Time-Discounted Ergodic Objective
Towards emphasizing the need to achieve ergodicity faster in time, in contrast to the infinite time needed for asymptotic ergodicity, a time-discounted ergodicity is defined next.
@@
[…]
@@ IV-D Symmetric Stochastic Matrix
The following shows that for symmetric stochastic matrices $P$, time-discounted ergodicity can be exactly optimized by minimizing the second largest eigenvalue of $P$. This is in contrast to minimizing the second largest eigenvalue modulus (SLEM) of $P$, which would speed up convergence to the desired stationary distribution, e.g., as in [15]. Specifically, to optimize the convergence rate for an arbitrary initial condition $\rho_{0}$, the stochastic matrix $P$ is selected to minimize the difference, normalized by the initial difference, between the expected time-discounted-average $\mathbb{E}[\hat{\rho}_{w}]$ and the space average $\bar{\rho}$ over different initial distributions $\rho_{0}$, i.e to minimize the normalized ergodicity deviation (NED)
$J_{\mathbb{E}}~{}=\max_{\rho_{0}}\frac{\left(\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right)^{T}Q\left(\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right)}{\left(\rho_{0}-\bar{\rho}\right)^{T}Q\left(\rho_{0}-\bar{\rho}\right)}$ (18)
$\qquad\qquad~{}=\max_{\rho_{0}}\frac{\|Q^{1/2}\left(\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right)\|_{2}}{\|Q^{1/2}\left(\rho_{0}-\bar{\rho}\right)\|_{2}}$
where $Q$ is a given weighting matrix.
The following theorem shows that for symmetric stochastic matrices $P$, under conditions typical for discount factors on the weighting scheme $w$, optimal time-discounted ergodicity (across all initial distributions) can be achieved by minimizing the second largest eigenvalue of $P$.
@@
Let the stochastic matrix $P$ be symmetric and irreducible. Moreover, let the time-discounted-average operator $f_{w}$ in (15) be positive semi-definite and strictly monotonic in $[-1,1]$ namely
$f_{w}(\lambda_{1})>f_{w}(\lambda_{2})\geq 0\quad\text{if}\quad\lambda_{1}>\lambda_{2},\quad\forall\lambda_{1},\lambda_{2}\in[-1,1].$ (19)
Then, the normalized ergodicity deviation (NED) in (18) with weighting matrix $Q=I$ is minimized, i.e.
$\min_{P}\left(\max_{\rho_{0}}\frac{\|\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\|_{2}}{\|\rho_{0}-\bar{\rho}\|_{2}}\right)$ (20)
for the stationary distribution $\bar{\rho}=\frac{1}{n}\textbf{1}$, if and only if the second largest eigenvalue (SLE) is minimized, i.e.,
@@
Since $P$ is a symmetric stochastic matrix, the eigenvalues of $P$, $\lambda(P)$, are real and lies within the domain $[-1,1]$, and it can be diagonalized as
$P$ $=V\Lambda U^{T}.$ (22)
where, $\Lambda$ is a diagonal matrix of all the eigenvalues of $P$, $V$ is a matrix with it’s column being the right eigenvectors of $P$, and $U$ is a matrix with it’s column being the left eigenvectors of $P$ ordered according to $\Lambda$.
Moreover, $f_{w}$ is a power series, and the set of eigenvalues $\lambda(P)$ lies within it’s domain, $f_{w}(P)$, obtained by replacing the scalar $\lambda$ in (15) with the matrix $P$,
@@
exists, has the same eigenvectors as $P$, and the eigenvalues of $f_{w}(P)$ are the scalar mapping $f_{w}(\lambda(P))$ of the eigenvalues $\lambda(P)$ of the matrix $P$ [43, Definition 1.2. on p.3], i.e.,
@@
$\left\|\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right\|_{2}$ (25)
$\quad=\left\|f_{w}(P)\rho_{0}-\bar{\rho}\right\|_{2}$
$\quad=\left\|Vf_{w}(\Lambda)U^{T}\rho_{0}-\bar{\rho}\right\|_{2}$
$\quad=\left\|f_{w}(\lambda_{1})v_{1}u_{1}^{T}\rho_{0}+\left(\sum_{i=2}^{n}f_{w}(\lambda_{i})v_{i}u_{i}^{T}\rho_{0}\right)-\bar{\rho}\right\|_{2}$
@@
$\left\|\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right\|_{2}$ (26)
$\quad=\left\|f_{w}(1)\bar{\rho}\textbf{1}^{T}\rho_{0}+\left(\sum_{i=2}^{n}f_{w}(\lambda_{i})v_{i}u_{i}^{T}\rho_{0}\right)-\bar{\rho}\right\|_{2}.$
Using $1\textbf{1}^{T}\rho_{0}=$ 1 for any probability-distribution vector, and $f_{w}(1)=1$ by construction in (15), the previous expression can be rewritten as
$\left\|\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right\|_{2}$ (27)
$\quad=\left\|\bar{\rho}+\left(\sum_{i=2}^{n}f_{w}(\lambda_{i})v_{i}u_{i}^{T}\rho_{0}\right)-\bar{\rho}\right\|_{2}$
$\quad=\left\|\sum_{i=2}^{n}f_{w}(\lambda_{i})v_{i}u_{i}^{T}\rho_{0}\right\|_{2}$
$\quad=\left\|V_{\perp}f_{w}(\Lambda_{\perp})U_{\perp}^{T}\rho_{0{\perp}}\right\|_{2}$
@@
Here, $P|_{u_{1}^{\perp}}$ is the restriction of $P$ in the subspace orthogonal to $u_{1}$, $\rho_{0{\perp}}$ is the component of $\rho_{0}$ that is orthogonal to $u_{1}$, and $V_{\perp}$,$I_{\perp}$, $\Lambda_{\perp}$ are the associated restrictions of $V$,$U$, $\Lambda$ in  (25), respectively. Then, the NED objective in (18) can be written using (27) as
@@
Since the set of $\rho_{0\perp}=\epsilon(\rho_{0})x$ spans the orthogonal space (see lemma in appendix A), the right hand side (rhs) of (28) is equivalent to the induced matrix norm as
@@
$=$ $\max_{\|x\|_{2}=1}\left\|f_{w}\left(P|_{u_{1}^{\perp}}\right)x\right\|_{2}$
$=$ $\left\|f_{w}\left(P|_{u_{1}^{\perp}}\right)\right\|_{2}.$
@@
where
$\lambda_{\text{max}}\left(P|_{u_{1}^{\perp}}\right)=\max_{i\neq 1}\lambda_{i}(P).$ (32)
From (28)-(32),
$\arg\min_{P}\left(\max_{\rho_{0}}\frac{\|\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\|_{2}}{\|\rho_{0}-\bar{\rho}\|_{2}}\right)$ $=\arg\min_{P}\left(\left\|f_{w}\left(P|_{u_{1}^{\perp}}\right)\right\|_{2}\right)$ (33)
$=\arg\min_{P}\left(\max_{i\neq 1}\lambda_{i}(P)\right)$
@@
The time-discounted-average operator $f_{w}$ in Eq. (15) does not satisfy the strictly monotonic condition in Theorem 2 with uniform weighting $w=(1,1,1,\cdots)$, since
$f_{w}(\lambda){}$ =$=\begin{cases}0&\text{ if }\lambda\in[-1,1)\\ 1&\text{ if }\lambda=1\end{cases}$ (34)
Therefore, for any symmetric and irreducible stochastic matrix $P$ satisfying $P\bar{\rho}=\bar{\rho}$ in (10) with $\bar{\rho}=\frac{1}{n}\textbf{1}$, the ergodic deviation becomes zero from (27) , i.e.,
@@
$\quad P_{i,j}=0\quad\text{if}\quad(j,i)\notin\mathcal{E}$ (Transitions $~{}~{}{(\mbox{Transitions in}}~{}\mathcal{G}).$
@@
The optimization follows since the second largest eigenvalue $\max_{i\neq 1}\lambda_{i}(P)$ in (32) is the largest eigenvalue of $P-\frac{2}{n}\textbf{1}\textbf{1}^{T}$, which is a Wielandt deflation [44] of the largest eigenvalue $\lambda_{1}$ of $P$ from $\lambda_{1}=1$ to $-1$ ( the smallest possible eigenvalue of a symmetric stochastic matrix), i.e.,
$\lambda(P-\frac{2}{n}\textbf{1}\textbf{1}^{T})$ =$=\{\lambda_{1}-2,\lambda_{2},\lambda_{3},\cdots\}$ (37)
=$=\{-1,\lambda_{2},\lambda_{3},\cdots\}$
@@
The minimization of the largest eigenvalue of a symmetric matrix ($P-\frac{2}{n}\textbf{1}\textbf{1}^{T}$) is convex [45, Example 3.10 p.82] and all the constraints are linear, and therefore  (36) is a convex optimization problem.

### v2

@@ Abstract
Ergodic exploration has spawned a lot of interest in mobile robotics due to its ability to design time trajectories that match desired spatial coverage statistics. However, current ergodic approaches are for continuous spaces, which require detailed sensory information at each point and can lead to fractal-like trajectories that cannot be tracked easily. This paper presents a new ergodic approach for graph-based discretization of continuous spaces. It also introduces a new time-discounted ergodicity metric, wherein early visitations of information-rich nodes are weighted more than late visitations. A Markov chain synthesized using a convex program is shown to converge more rapidly to time-discounted ergodicity than the traditional fastest mixing Markov chain. The resultant ergodic traversal method is used within a hierarchical framework for active inspection of confined spaces with the goal of detecting anomalies robustly using SLAM-driven Bayesian hypothesis testing. Experiments on a ground robot show the advantages of this framework over three continuous space ergodic planners as well as greedy and random exploration methods for left-behind foreign object debris detection in a ballast tank.
@@ I Introduction
There are three main challenges of continuous-space ergodic control for inspection of confined spaces, all of which are addressed in our discrete-space method. First, the continuous space methods tend to yield complex, fractal-like trajectories to satisfy the ergodicity requirements, which are difficult to execute in confined spaces. Moreover, the complex trajectories (i) can potentially destabilize robot localization by introducing large nonlinearities in the inputs and (ii) can cause the robot to inspect local featureless areas for long periods of time. A graph-based discretization of the confined space into relatively large regions allows designing a region-wise ergodic traversal method that works without considering local collision-free motion planning directly. Second, the information field for the continuous-space ergodic method requires a sensor model to consider occlusion at every point in space. In contrast, the information is bulk estimated at each node on the graph (region in the space) in our discrete ergodic control method. Third, the continuous-space ergodic methods cannot model directional traversability constraints in the confined space, such as an unclimbable slope in one direction, or a door that only opens on one side. In contrast, such connectivity constraints are modeled as directed graph edges in our discrete method
@@
This graph-based discretization enables the development of a hierarchical planning framework, which allows decomposition and modularization of both robot control and estimation to reduce the overall computational complexity of active inspection. We focus on anomaly detection as the inspection goal. Consequently, REMC sits at the top of the hierarchy as the region-level planner, where coarse-grained information on the current uncertainty estimates of region-wise anomalies are fed from the region graph. At the second level, fine-grained information on anomaly uncertainty estimates within the regions are used for inspection waypoint placement. At the bottom level, robot state estimates and an occupancy grid map are provided as inputs by a graph-SLAM algorithm for local waypoint navigation. The uncertainty estimates come from a Bayesian anomaly detector at this level, which recursively updates the posterior of the robot’s local observations corresponding to anomalies based on the signed Mahalanobis distances of the observations from their nearest points on a reference model of the confined space. The overall inspection framework is termed hierarchical ergodic Markov planner (HEMaP). Both simulation and physical experiments are carried out inside a representative ballast tank with FODs as anomalies, using a ground robot equipped with a depth camera and GPU-based on-board processing. The results show that HEMaP achieves high FOD detection rates (averaging 85%), and consistently outperforms greedy and random exploration methods regardless of the allowable inspection time frame.
@@ II-A Ergodic Control: Continuous Versus Discretized Space
There are three main branches of work that attempt to address these challenges. One branch utilizes the Fourier metric as an objective function and expands the method into a trajectory optimization problem, where collision avoidance is handled as an optimization constraint [16, 17, 18, 19]. However, it inherits a similar problem as the Fourier-based metric, where the finite number of frequency coefficients are not able to capture complex obstacle geometry such as structures in the confined space. When modeled as constraints, the complex geometries, together with the robot motion model, introduce high non-convexity, which generates more computational complexity. The second branch of work uses the radial basis function instead of the Fourier basis to adapt to different domain topologies [20, 21, 22, 23]. The basis functions are specific to the domain topology and are pre-computed using prior maps. When unexpected obstacles such as FODs appear, a new set of basis functions may be required. In contrast, new obstacles are handled by a standard local planner in our hierarchical approach. The third line of work changes the ergodic metric to KL-divergence between the robot distribution and the target distribution. As a result, this metric is able to adapt to a wider range of sensor models and has a lower computational complexity than the Fourier method [24]. This line of work is similar to our approach in that both utilize stochasticity to solve the ergodic control problem. However, instead of using stochastic optimization to efficiently sample trajectories that solve the finite-horizon optimization problem, we find a stochastic policy that is inherently ergodic.
Several works have adapted these three branches of ergodic control methods on various robotics tasks, including target localization [25], multi-link manipulator control [26, 27, 28, 29], and swarm control [30]. In these tasks, the information field is either given or has a well-behaved analytical form. However, information is harder to model in inspection applications, which instead benefit from a graph-based abstraction. Recently, advances have also been made on ergodic exploration of cluttered environments with many obstacles. For example, this has been addressed by sampling random graphs in the free space and then using a Dijkstra-like method to search for an ergodic path [31]. However, this method is computationally challenging and does not guarantee convergence. Alternatively, the method presented in [20] has been adapted to a maze-like environment [32]. However, this is only applicable to square-lattice graphs instead of arbitrary strongly connected graphs, as in our method. Ergodicity in continuous space with arbitrary topology has also been achieved by a Markov process that satisfies measure-preserving flows [33]. A similar Markov process is used in our method, but over graphs in discretized spatial regions, which formulates a linear specification of the transition model and enables convex optimization of the convergence rate to ergodicity.
@@ III-A Hierarchical Representation of Workspace
We consider the set of regions in the workspace as an arbitrary coarse graining of the pose space $\mathcal{X}\subset\text{SE}(3)$ into a finite disjoint set $\mathcal{R}=\{r_{1},r_{2},\cdots,r_{n}\}$, where the space within each region is connected.
@@ IV-A Continuous and Discrete Ergodic System
for all initial conditions $r[0]$. Here, the right-hand side is the space average of $F$ over all elements $r$ in the set $\mathcal{R}$ with respect to a measure $\mu$ over the sigma algebra $\Sigma$. The left-hand side is the (long-term) time average of $F$ evaluated along the sequence $r[k]$ generated by a dynamic map $r[k+1]=M(r[0:k]):\mathcal{R}\rightarrow\mathcal{R}$ with initial state $r[0]$.
The graph-based ergodicity in (IV.2) implies that the relative frequency of the robot visiting each region (i.e., the time average $\hat{\rho}$) is equal to the relative measure of each region on the graph (i.e., the space average $\bar{\rho}$) if the measurable function $F$ in (3) is selected as the indicator function II
@@ IV-B Asymptotic Ergodicity
We propose to achieve graph-based ergodicity by using a Markov chain, where the transition probability between regions is determined by a stochastic matrix $P$ with the stationary distribution equal to the desired distribution. Then, at the limit, the space average is almost surely equal to the desired distribution based on the ergodic theorem of Markov chains [14, Theorem 1.10.2].
Specifically, consider a stochastic dynamic map $M$ , where the transition probability $P_{j,i}$ of the robot from region $r_{i}$ to $r_{j}$ forms a stochastic matrix $P\in\mathbb{R}^{n\times n}$. Then, asymptotic ergodicity, in the expected sense, is achieved for any stochastic matrix $P$ that is irreducible, i.e., can go from any region to any other region in the graph $\mathcal{G}$ (not necessarily in one move), with the space average $\bar{\rho}$ as the unique stationary distribution, as shown below.
Theorem 1 (Ergodic Theorem for Ergodic Graph Traversal).
@@
Remark 1.
The term ergodic Markov chain has multiple definitions in the literature, e.g., when the Markov chain is irreducible [46, p.245], meaning it can transition from any state to any other state in a finite amount of time; or when the Markov chain is both irreducible and aperiodic [47, p.316]. In this paper, we adopt the definition most closely aligned with ergodic theory [14, Theorem 4.1], which requires irreducibility (but not aperiodicity) and further stipulates that the time average must equal a specific space average as in Definition (IV.2).
@@ IV-C Time-Discounted Ergodic Objective
In contrast to the infinite time for asymptotic ergodicity, a time-discounted ergodicity is defined next to formalize the need for faster ergodicity in time.
@@
$\begin{aligned} f_{w}(\lambda)\triangleq\lim_{K\to\infty}\frac{1}{\sum_{k{=0}}^{K-1}1/k!}\left(\sum_{k=0}^{K-1}\frac{1}{k!}\lambda^{k}\right)\end{aligned}~=e^{-1}{e^{\lambda}}.$ (15)
In general, discounted ergodicity (11) cannot be achieved for an arbitrary initial condition. Instead, the objective is to select the stochastic matrix $P$ optimally to reduce the difference between the expected time-discounted average $\mathbb{E}[\hat{\rho}_{w}]$ in Eq. (12) and the space average $\bar{\rho}$ over different initial distributions $\rho_{0}$, normalized over the initial difference. In other words, we want to minimize the normalized ergodicity deviation (NED)
$J_{\mathbb{E}}~=\max_{\rho_{0}}\frac{\|Q^{1/2}\left(\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right)\|_{2}}{\|Q^{1/2}\left(\rho_{0}-\bar{\rho}\right)\|_{2}},$ (16)
where $Q$ is a given weighting matrix [^1] A popular method for defining the convergence rate is to calculate the error for an arbitrarily finite time horizon. This is a special case of the discounted time average with a weighting function $w=[1,1,1\cdots 0,0,0\cdots]$..
@@ IV-D Symmetric Stochastic Matrix
The following shows that for symmetric stochastic matrices $P$, time-discounted ergodicity can be exactly optimized by minimizing the second largest eigenvalue of $P$. This is in contrast to minimizing the second largest eigenvalue modulus (SLEM) of $P$, which would speed up convergence to the desired stationary distribution, e.g., as in [15].
The theorem below shows that for symmetric stochastic matrices $P$, under conditions typical for discount factors on the weighting scheme $w$, optimal time-discounted ergodicity (across all initial distributions) can be achieved by minimizing the second largest eigenvalue of $P$.
@@
Let the stochastic matrix $P$ be symmetric and irreducible. Moreover, let the time-discounted-average operator $f_{w}$ in (13) be positive semi-definite and strictly monotonic in [1,1][-1,1] namely $f_{w}(\lambda_{1})>f_{w}(\lambda_{2})\geq 0$ if $\lambda_{1}>\lambda_{2}$ for all $\lambda_{1},\lambda_{2}\in[-1,1]$. Then, the normalized ergodicity deviation in (16) with weighting matrix $Q=I$ is minimized for the stationary distribution $\bar{\rho}=\frac{1}{n}\textbf{1}$, if and only if the second largest eigenvalue (SLE) is minimized, i.e.,
@@
Since $P$ is a symmetric stochastic matrix, the eigenvalues of $P$, $\lambda(P)$, are real and lies within the domain [1,1][-1,1], and it can be diagonalized as $P=V\Lambda U^{T}$, where $\Lambda$ is a diagonal matrix of all the eigenvalues of $P$, $V$ is a matrix with its column being the right eigenvectors of $P$, and $U$ is a matrix with its column being the left eigenvectors of $P$ ordered according to $\Lambda$. Moreover, $f_{w}$ is a power series, and the set of eigenvalues $\lambda(P)$ lies within it’s domain, $f_{w}(P)$, obtained by replacing the scalar $\lambda$ in (13) with the matrix $P$,
@@
exists, has the same eigenvectors as $P$, and the eigenvalues of $f_{w}(P)$ are the scalar mapping $f_{w}(\lambda(P))$ of the eigenvalues $\lambda(P)$ of the matrix $P$ [48, Definition 1.2. on p.3], i.e.,
@@
$\left\|\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right\|_{2}=\left\|f_{w}(P)\rho_{0}-\bar{\rho}\right\|_{2}=\left\|Vf_{w}(\Lambda)U^{T}\rho_{0}-\bar{\rho}\right\|_{2}$ (20)
$~=\left\|f_{w}(\lambda_{1})v_{1}u_{1}^{T}\rho_{0}+\left(\sum_{i=2}^{n}f_{w}(\lambda_{i})v_{i}u_{i}^{T}\rho_{0}\right)-\bar{\rho}\right\|_{2}$
@@
$\left\|\mathbb{E}[\hat{\rho}_{w}]\!-\bar{\rho}\right\|_{2}\!=\!\!\left\|f_{w}(1)\bar{\rho}\textbf{1}^{T}\rho_{0}+\left(\sum_{i=2}^{n}f_{w}(\lambda_{i})v_{i}u_{i}^{T}\rho_{0}\right)-\bar{\rho}\right\|_{2}\!.$ (21)
Using $\textbf{1}^{T}\rho_{0}=1$ for any probability-distribution vector, and $f_{w}(1)=1$ by construction in (13), the previous expression can be rewritten as
$\left\|\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho}\right\|_{2}=\left\|\bar{\rho}+\left(\sum_{i=2}^{n}f_{w}(\lambda_{i})v_{i}u_{i}^{T}\rho_{0}\right)-\bar{\rho}\right\|_{2}$ (22)
$\quad=\left\|\sum_{i=2}^{n}f_{w}(\lambda_{i})v_{i}u_{i}^{T}\rho_{0}\right\|_{2}~=\left\|V_{\perp}f_{w}(\Lambda_{\perp})U_{\perp}^{T}\rho_{0{\perp}}\right\|_{2}$
@@
Here, $P|_{u_{1}^{\perp}}$ is the restriction of $P$ in the subspace orthogonal to $u_{1}$, $\rho_{0{\perp}}$ is the component of $\rho_{0}$ that is orthogonal to $u_{1}$, and $V_{\perp}$$I_{\perp}$, $\Lambda_{\perp}$ are the associated restrictions of $V$,$U$, $\Lambda$ in  (20), respectively. Then, the NED objective in (16) can be written using (22) as
@@
Since the set of $\rho_{0\perp}=\epsilon(\rho_{0})x$ spans the orthogonal space, (see lemma in Appendix A), the right hand side (rhs) of (23) is equivalent to the induced matrix norm as
@@
$\quad=\max_{\|x\|_{2}=1}\left\|f_{w}\left(P|_{u_{1}^{\perp}}\right)x\right\|_{2}=\left\|f_{w}\left(P|_{u_{1}^{\perp}}\right)\right\|_{2}.$
@@
where $\lambda_{\text{max}}\left(P|_{u_{1}^{\perp}}\right)=\max_{i\neq 1}\lambda_{i}(P)$. From (23)-(26),
$\arg\min_{P}\left(\left\|f_{w}\left(P|_{u_{1}^{\perp}}\right)\right\|_{2}\right)=\arg\min_{P}\left(\max_{i\neq 1}\lambda_{i}(P)\right)$ (27)
@@
The time-discounted-average operator $f_{w}$ in Eq. (13) does not satisfy the strictly monotonic condition in Theorem 2 with uniform weighting $w=(1,1,1,\cdots)$, since $f_{w}(\lambda)=0$ if $\lambda\in[-1,1)$ and $f_{w}(\lambda)=1$ $\lambda=1$. Therefore, for any symmetric and irreducible stochastic matrix $P$ satisfying $P\bar{\rho}=\bar{\rho}$ in (10) with $\bar{\rho}=\frac{1}{n}\textbf{1}$, the ergodic deviation becomes zero from (22) , i.e.,
@@
$\quad P_{i,j}=0\quad\text{if}\quad(j,i)\notin\mathcal{E}$ $~~{(\mbox{Transitions in}}~\mathcal{G}).$
@@
The optimization follows since the second largest eigenvalue $\max_{i\neq 1}\lambda_{i}(P)$ is the largest eigenvalue of $P-\frac{2}{n}\textbf{1}\textbf{1}^{T}$, which is a Wielandt deflation [49] of the largest eigenvalue $\lambda_{1}$ of $P$ from $\lambda_{1}=1$ to 1-1 ( the smallest possible eigenvalue of a symmetric stochastic matrix), i.e.,
$\lambda(P-\frac{2}{n}\textbf{1}\textbf{1}^{T})$ $=\{\lambda_{1}-2,\lambda_{2},\lambda_{3},\cdots\}$ (30)
$=\{-1,\lambda_{2},\lambda_{3},\cdots\}$
@@
The minimization of the largest eigenvalue of a symmetric matrix ($P-\frac{2}{n}\textbf{1}\textbf{1}^{T}$) is convex [14, Example 3.10 p.82] and all the constraints are linear, and therefore  (29) is a convex optimization problem. ∎